# What We Talk About When We Talk About LMs: Implicit Paradigm Shifts and the Ship of Language Models

### v2

@@ 5 Concluding Remarks
Our findings depict the most dramatic reshuffles of models, while the nuances of LMs can be found similarly at the underlying architecture level, or different variations of models. Besides, causal links between our work and the identified key factors in prior works may help depict the field’s trajectories, and qualitative studies such as interviews with paper authors can also help understand the different patterns and beliefs. Beyond the LM-specific discussions, we also highlight the broad existence of the Ship of Theseus scenario in various other terms, topics, and fields.
@@
While our methodologies can be naturally extended, we would like to note that the current analyses and implementations have limitations. First, although we have gone over various findings in recent years with the most dramatic advances of LMs, a holistic overview of a research topic, let alone an entire field, is not covered by the scope of a few years. It would be meaningful if these most rapid changes could be connected with the decades of conceptualization and exploration preceding the engineering breakthroughs. Similarly, using main conference papers at the major international venues as a proxy of the NLP community has its limitations. We encourage future works to take broader consideration of the essential contributions that are less represented by the relatively convergent selections of such venues, e.g., regional conferences on dialect or indigenous language.
@@
Our data is collected from the ACL Anthology on the terms of Creative Commons 4.0 BY (Attribution) license, which allows unlimited reproduction, distribution, and hosting of materials for non-commercial purposes [^6] See CC BY 4.0 and the Anthology’s copyright statement for more information.. The authors report no other potential ethical considerations.
@@ Appendix A The dataset construction pipeline
Figure 8: The full pipeline for constructing the model dictionaries. The LLM agent follows a formatted prompt (§A.4) to automatically identify potential models. The extracted strings are merged and ranked by frequency to form the list of candidate names. Then, we manually determine whether it is a new entry, an alias, or other based on (1) whether the candidate is indeed a valid model name, (2) whether it refers to the same model as an existing entry does, and (3) whether it’s already covered by an alias.
We provide more details about the semi-automated dataset creation process, including the illustration of the workflow, the full list of models involved, and the prompts used for the automated part of model name identification. The full dataset and more details will be published for future use in the final version of this project.
@@ A.1 A note on the scope of “LMs”
To accommodate the underspecification and the even more vague real-world usages, we thus consider the “broadest” definitions that are historically involved. This would include text embedding schemes that have essentially been used in the place of “LMs”, key parts of an LM-like system where the system is not otherwise named, or artifacts created in the “LM” way but are not exactly a “model”. (Note that this work concerns “what we talk about when we talk about LMs”, not just “which LM we talk about when we talk about LMs”.)
This is especially worth noting since the terminology of “LM” is not always the most prevalent. As an important instance, “word embeddings/vectors” had once been a heated topic; it is both true that (1) the terms are not equivalent to “LMs” strictly speaking, and that (2) it was a very common (if not default) practice to build (e.g. add a linear layer) on top of the embeddings to obtain probabilities/likelihood of the text in question. This is essentially close to “LM” as defined, though we rarely referred to the entire system this way but would have often used wordings like “word vectors + [classification/prediction/…]” back then. To this end, we do consider these “broader” senses because they have long been used for the same sense and in the same context. In fact, our motivations align with such situations: the communities’ description and perception of a concept might have undergone many changes and variations, but we are yet to know what exactly has changed, to what extent, and how varied it is.
@@
A.2 Classifying potential model names
See Figure 8.
A.3 Full list of models
ChatGPT, GPT-3, GPT-4, BERT, T5, GPT-3.5, GPT-2, LLaMA, RoBERTa, PaLM, CLIP, BART, XLM-R, Alpaca, BLOOM, mT5, InstructGPT, mBERT, GPT-J, Flan-T5, OPT, Codex, COMET, ELECTRA, Longformer, mBART, SimCSE, BLOOMZ, BigBird, BLIP, DeBERTa, CodeT5, Switch Transformer, Vicuna, T0, PEGASUS, LSTM, ALBERT, DPR, Macaw, LXMERT, SpanBERT, TinyBERT, ViLBERT, XLM, Linformer, kNN-LM, kNN-MT, REALM, RETRO, GraphCodeBERT, Sentence-BERT, RNN, HyperCLOVA, CodeGen, Dolly, Pythia, LaMDA, FLAN, BLIP-2, XLNet, GPT, ELMo, BioBERT, DialoGPT, RemBERT, PaLM 2, DistilBERT, SciBERT, ClinicalBERT, M2M100, GloVe, LASER, word2vec, fastText, LaBSE, CNN, wav2vec, UNITER, MASS, MT-DNN, BlenderBot, DistMult, OFA, CMLM, HRED, ERNIE, ConveRT, MiniLM, Galactica, RuleTakers, Claude, LayoutLM, ST-DNN, IRNet
@@ A.4 Prompts for automated model name extraction
Appendix B Sunburst graphs
The following pages show the Sunburst graphs for all 10 conferences in chronological order.

### v3

@@ 5 Concluding Remarks
Our findings depict the most dramatic reshuffles of models, while the nuances of LMs can be found similarly at the underlying architecture level, or different variations of models. Besides, causal links between our work and the identified key factors in prior works may help depict the field’s trajectories, and qualitative studies such as interviews with paper authors can also help understand the different patterns and beliefs. Beyond the LM-specific discussions, we also highlight the broad existence of the Ship of Theseus scenario in various other terms, topics, and fields. For instance, it would be intriguing if similar analyses can be applied to understand how the interpretations of fairness or safety in the AI context were enriched over time.
@@
While our methodologies can be naturally extended, we would like to note that the current analyses and implementations have limitations. Although we reviewed various findings in recent years with the most dramatic advances of LMs, a holistic overview of a research topic, let alone an entire field, is not covered by the scope of a few years. It would be meaningful if these most rapid changes could be connected with the decades of conceptualization and exploration preceding the engineering breakthroughs. Similarly, using main conference papers at the major international venues as a proxy of the NLP community has its limitations. We encourage future works to take broader consideration of the essential contributions that are less represented by the relatively convergent selections of such venues, e.g., regional conferences on dialect or indigenous language.
@@
Our data is collected from the ACL Anthology on the terms of Creative Commons 4.0 BY (Attribution) license, which allows unlimited reproduction, distribution, and hosting of materials for non-commercial purposes [^7] See CC BY 4.0 and the ACL Anthology’s copyright statement for more information.. The authors report no other potential ethical considerations.
@@ Appendix A The dataset construction pipeline
We provide more details about the semi-automated dataset creation process, including the full list of model names, and the prompts and parameters for the automated part of model name identification.
@@ A.1 A note on the scope of “LMs”
To accommodate the underspecification and the even more vague real-world usages, we thus consider the “broadest” definitions that are historically involved. This would include text embedding schemes that have essentially been used in the place of “LMs”, key parts of an LM-like system where the system itself is not otherwise named, or artifacts created in the “LM” way but are not exactly a “model”. (Note that this work concerns “what we talk about when we talk about LMs”, not just “which LM(s) we talk about when we talk about LMs”.)
This is especially worth noting since the terminology of “LM” is not always the most prevalent. As an important instance, “word embeddings/vectors” had been a central topic; it is both true that (1) the terms are not equivalent to “LMs” strictly speaking, and that (2) it was a very common (if not the default) practice to build (e.g. add a linear layer) on top of the embeddings to obtain probabilities/likelihood of the text in question. This is essentially close to “LM” as defined, though we rarely referred to the entire system this way but would have often used wordings like “word vectors + [classification/prediction/…]” back then. To this end, we do consider these “broader” senses because they have long been used for the same sense and in the same context. In fact, our motivations align with such situations: the communities’ description and perception of a concept might have undergone many changes and variations, but we are yet to know what exactly has changed, to what extent, and how varied it is.
@@
A.2 Full list of models
ChatGPT, GPT-3, GPT-4, BERT, T5, GPT-3.5, GPT-2, LLaMA, RoBERTa, PaLM, CLIP, BART, XLM-R, Alpaca, BLOOM, mT5, InstructGPT, mBERT, GPT-J, Flan-T5, OPT, Codex, COMET, ELECTRA, Longformer, mBART, SimCSE, BLOOMZ, BigBird, BLIP, DeBERTa, CodeT5, Switch Transformer, Vicuna, T0, PEGASUS, LSTM, ALBERT, DPR, Macaw, LXMERT, SpanBERT, TinyBERT, ViLBERT, XLM, Linformer, kNN-LM, kNN-MT, REALM, RETRO, GraphCodeBERT, Sentence-BERT, RNN, HyperCLOVA, CodeGen, Dolly, Pythia, LaMDA, FLAN, BLIP-2, XLNet, GPT, ELMo, BioBERT, DialoGPT, RemBERT, PaLM 2, DistilBERT, SciBERT, ClinicalBERT, M2M100, GloVe, LASER, word2vec, fastText, LaBSE, CNN, wav2vec, UNITER, WizardLM, MASS, MT-DNN, BlenderBot, OFA, CMLM, HRED, ERNIE, ConveRT, Diffusion-LM, MiniLM, Falcon, Galactica, PPLM, RuleTakers, Claude, Tk-Instruct, LayoutLM, PanGu-$\alpha$, GROVER, CTRL, EntityNLM, ST-DNN, SpERT
A.3 Settings of the automated model name detection module
The LLM-based name detection is performed with OpenAI’s gpt-4-0125-preview (also known as gpt-4-turbo-preview) [^8] https://platform.openai.com/docs/models#gpt-4-turbo-and-gpt-4, accessed in February to June 2024. The temperature is set to 0, and we follow the default settings of other parameters, e.g., top_p = 1.00.
@@ A.4 Prompts for automated model name extraction
Appendix B Sunburst diagrams
The following pages show the Sunburst diagrams for all 10 conferences in chronological order.